two guns are pointed at each other one upward at 30° and other at same angle of depression. The nozzle being 30m apart. Of the guns are shot with a velocity of 350 m/s upward and 300m/s downward. Find when and where they will meet.


1
Expert's answer
2019-09-10T13:23:23-0400

We have:


"x_1+x_2=x"

"v_1\cos{30\degree}t+v_1\cos{30\degree}t=x"

"t=\frac{30}{(350+300)cos{30\degree}}=0.053\ s"

The distance from first gun:


"d=x_1=v_1\cos{30\degree}t=(350)(0.053)\cos{30\degree}=16\ m"